OpenStudy (anonymous):

HELP will give medal and become fan!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! In many localities firewood is sold by the cord, which is a pile 4 feetwide, 4feet high, and 8 feet long, containing 128 cubic feet. If the width and the height of the pile remain constant, the number of cords (c) of wood varies directly as the length of the pile. If a pile of wood 12 feet long contains1.5 cords how many will there be in a pile 10 feet long? 16 feet long? 30 feet long? How long would the pile be to connect 5.5 cords? 3.75?

OpenStudy (anonymous):

W = width H = height your 12 feet long 1.5 cord pile: \[1.5 \times 128 = 12 \times W \times H\] \[WH = 16\] Volume = LWH plug in the different Ls since WH is constant. the amount of cords will likely be a fraction:\[\frac{ answer }{ 128 }= cords\]